OpenStudy (anonymous):

tan theta=3/8 third quadrant,find cos theta

OpenStudy (anonymous):

 \[\sec ^{2}\theta-\tan ^{2}\theta=1,\sec ^{2}\theta=1+\tan ^{2}\theta=1+\frac{ 9 }{ 64 }=\frac{ 73 }{64 }\] \[\cos ^{2}\theta=\frac{ 64 }{73 },\cos \theta=\frac{- 8 }{\sqrt{73} }\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw